OpenStudy (anonymous):

find the vertex, line of symmetry and max/min value of f(x)=x^2-12x-6

OpenStudy (turingtest):

f'(x)=2x-12=0 for max's and min's x=6 is the min since the parabola opens upwards

OpenStudy (turingtest):

the line of symmetry is at the min so y=f(6)=36-36-6=-6 is the line